In a normal distribution, 8% of items are under 50 and 10% are over 60. Find the mean and standard deviation of the distribution. 


1
Expert's answer
2021-05-07T11:47:16-0400

Value of Z, corresponding to 0.5-0.08=0.42 are , is equal to 0.202


     "0.202=\dfrac{50-\mu}{\sigma}"


      "\Rightarrow \mu-0.202\sigma=50~~~~~~~~~-(1)"


Value of Z, corresponding to 0.5-0.10=0.40 are , is equal to 0.253


      "0.253=\dfrac{60-\mu}{\sigma}"


    "\Rightarrow \mu-0.253\sigma=60~~~~~~~~~-(2)"


Solving equation 1 and 2 we get-

"\mu=10.39\\

 \sigma=2.53"